RhondaSommer (rhondasommer):

An airplane is flying 4,200 feet above the ground at an angle of depression of 7.8°. How far must the plane fly in order to land? Round your answer to the nearest foot. 4,239 feet 30,947 feet 30,661 feet 32,583 feet

OpenStudy (isaidavila):

Ctg(7.8)=x/4200 7.304=x/4200 Solve that and aproximate

RhondaSommer (rhondasommer):

thanks

RhondaSommer (rhondasommer):

so... C right? @isaidavila

OpenStudy (isaidavila):

Aproximately